OpenStudy (anonymous):

how many atoms are in .20 moles of lead

OpenStudy (abb0t):

0.2 moles \(\times\) \(\frac{atom}{6.022 \times 10^{23}mol}\) = atoms of Lead

OpenStudy (anonymous):

@abb0t SorRy didn't gt it..

OpenStudy (abb0t):

Don't get what?

OpenStudy (anonymous):

0.2 moles \(\times\) \(\frac{atom}{6.022 \times 10^{23}mol}\) = atoms of Lead

OpenStudy (anonymous):

???

OpenStudy (anonymous):

hmm....

OpenStudy (theeric):

\(1\ mol\ of\ atoms\approx 6.022\times 10^{23}\ atoms\\\implies1=\dfrac{6.022\times 10^{23}\ atoms}{1\ mol\ of\ atoms}\) @abbot , isn't that the conversion factor? So if you have \(0.20\ mol\ of\ lead\ atoms\), you multiply it by \(1\), the conversion factor. \(0.20\ mol\ of\ lead\ atoms\times\dfrac{6.022\times 10^{23}\ atoms}{1\ mol\ of\ atoms}\huge\longrightarrow\normalsize atoms\ of\ lead\) @Khiz , is this what you were thinking?

OpenStudy (ookawaiioo):

@theEric is right. abbot and khiz got them mixed around.

OpenStudy (theeric):

Alright! I think @KhiZ was confused by @abbot 's response, and posted abbot's equation as part of a question.

OpenStudy (theeric):

Thanks for checking, too, @oOKawaiiOo ! :)

OpenStudy (ookawaiioo):

np!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Thanx @theEric :) Thatx exactly wht i ws thnking..

OpenStudy (theeric):

Cool! :) Glad we got that settled, then. :)

OpenStudy (anonymous):

yupx :)